Given y= 50+7x, answer the word problem:


  1. Apart from the basic charge of PHP50.00 and PHP7.00 per kilometer, the courier service also charges an additional PHP80.00 during afterhours (10pm to 6am) and waiting time charge of PHP60.00 per hour which is applicable 30 minutes after pickup or delivery time. How much will you pay the courier service if he will deliver fireworks from your home to Kilometer Zero in Luneta on February 11, 2021 at 10:30pm but he must wait until it will be lighted at 12 midnight?


Given y=2000/x


  1. Given that the van rental is PHP2000.00 per day and the gasoline that will be filled in the tank is 42 liters every week except on the third week where they only filled the van with 21 liters of gasoline that costs PHP45.00 per liter. How much will each of the 15 passengers (not including the driver) share in a month? (Consider that there are 4 weeks in a month and the passengers will travel 5 days a week.)

Thank you so much!


1
Expert's answer
2021-03-29T14:49:11-0400

1. y = 50 + 7x + 80 + 60 = 190 + 7x, where x is unknown distance from your home.

2. "y = (2000\u00d720 + 45\u00d7(42\u00d73 + 21))\/15 = 3,107.67."